OpenStudy (precal):

A solid is formed by adjoining two hemispheres to the ends of a right circular cylinder. The total volume of the solid is 12 cubic centimeters. Find the radius of the cylinder that produces the minimum surface area.

OpenStudy (precal):

I need to double check my set up.

OpenStudy (precal):

I used the volume of a sphere + volume of a cylinder

OpenStudy (precal):

\[V=\frac{ 4 }{ 3 }\pi r^3 + \pi r^2h\]

OpenStudy (precal):

I also, used the Surface Area formulas \[S=4\pi r^2+2\pi r h\]

OpenStudy (precal):

I set V = 12 and solved for h, then sub h into Surface Area formula

OpenStudy (precal):

I took the derivative of S and set it equal to zero, this is where I probably messed up because I got 9 pi as the radius

ganeshie8 (ganeshie8):

setup looks good to me

OpenStudy (precal):

9 pi is not a min value

OpenStudy (precal):

should I show my work so someone can point out my mistake, I would appreciate it

ganeshie8 (ganeshie8):

ohk.. whats the correct answer ?

OpenStudy (precal):

I don't know, this one I have no key to at all

OpenStudy (anonymous):

Yes, please show your work

OpenStudy (precal):

ok so V=12 so therefore \[12=\frac{ 4 }{ 3 }\pi r^3+\pi r^2h\]

OpenStudy (precal):

solving for h \[\frac{ 12-\frac{ 4 }{ 3 }\pi r ^3 }{ \pi r^2 }=h\]

OpenStudy (precal):

now I am going to sub h into the Surface Area formula\[S=4 \pi r^2+2 \pi r \left( \frac{ 12-\frac{ 4 }{ 3 }\pi r^3 }{ \pi r^2 } \right)\]

OpenStudy (precal):

see any mistakes so far?

ganeshie8 (ganeshie8):

nope, everything fine so far

OpenStudy (precal):

simplifying it a bit before I take the derivative of it is: \[S=4\pi r^2 +24r ^{-1}-\frac{ 8 }{ 3 }\pi r^2\]

OpenStudy (precal):

\[S=\frac{ 4 }{ 3 }\pi r^2+24r ^{-1}\]

OpenStudy (precal):

now I will take the derivative, see any mistakes?

OpenStudy (precal):

\[S ' =\frac{ 8 }{ 3 }\pi r - \frac{ 24 }{ r^2 }\]

OpenStudy (precal):

I set S'=0 then I solved for r

ganeshie8 (ganeshie8):

looks good, multiply 3r^2 and get rid of fractions

ganeshie8 (ganeshie8):

you must get cuberoots in ur answer

OpenStudy (precal):

ok did not think to take that approach

OpenStudy (precal):

need a moment to do it on paper

OpenStudy (precal):

so its the cube root of (9/pi)?

OpenStudy (precal):

now that is the min value I was looking for, thanks. I knew I had made a mistake solving for r....

ganeshie8 (ganeshie8):

\(\large r = \sqrt[3]{9/\pi}\) minimizes the surface area of solid

OpenStudy (precal):

thank you so much....

ganeshie8 (ganeshie8):

np :)